# Variable Neighborhood Searching Rerandomization

### v1

@@ Abstract
Rerandomization discards undesired treatment assignments to ensure covariate balance in randomized experiments. However, rerandomization based on acceptance-rejection sampling is computationally inefficient, especially when numerous independent assignments are required to perform randomization-based statistical inference. Existing acceleration methods are suboptimal and are not applicable in structured experiments, including stratified experiments and experiments with clusters. Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to draw balanced assignments in various experiments efficiently. We derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR. Simulation studies and a real data example indicate that our method maintains the appealing statistical properties of rerandomization and can sample thousands of treatment assignments within seconds, even in cases where existing methods require an hour to complete the task.
Keywords causal inference  $\cdot$ clinical trials  $\cdot$ experimental design  $\cdot$ combinatorial optimization  $\cdot$ binary quadratic programming  $\cdot$ variable neighborhood search
@@ 1 Introduction
Randomized experiments balance covariates on average and have been the gold standard for drawing causal inferences. However, as pointed out by Morgan and Rubin (2012), covariate imbalances between treatment groups often exist in practice due to the randomness in complete randomization(CR). An intuitive and practical solution to this issue is randomizing repeatedly until a prespecified covariate balance criterion is achieved. Despite the long history and widespread use of rerandomization (Student, 1938; Cox, 1982; Bailey and Rowley, 1987; Maclure et al., 2006; Imai et al., 2008; Bruhn and McKenzie, 2009), its theoretical implications were first studied by Morgan and Rubin (2012) using the Mahalanobis distance only recently. Since then, rerandomization has attracted growing research interests, and its theory underpinnings have been established in various scenarios (Morgan and Rubin, 2015; Li et al., 2018; Zhou et al., 2018; Li et al., 2020; Shi et al., 2022; Wang et al., 2023; Lu et al., 2023).
As the name of rerandomization indicates, researchers typically conduct rerandomization by repeatedly sampling treatment assignments until a balanced one is generated. This procedure is then repeated thousands of times since numerous independent qualified assignments are required to perform randomization-based inference. Therefore, rerandomization can be time-consuming using the acceptance-rejection sampling rerandomization(ARSRR) strategy. Zhu and Liu (2022) proposed a pair-switching rerandomization(PSRR) method to reduce the computational cost of rerandomization. However, the authors only considered a fixed number of covariates and a fixed covariate imbalance threshold. As we will show in Section 5, PSRR can be slower than ARSRR in certain settings. Moreover, both ARSRR and PSRR are computationally inefficient in settings considered by recent works, including diminishing covariate imbalance and high-dimensional covariates (Branson and Shao, 2021; Wang and Li, 2022; Zhang et al., 2023). Additionally, although Zhu and Liu (2022) extended PSRR to sequential experiments, PSRR is not applicable in other structured experiments such as stratified experiments or experiments with clusters.
@@
In this work, we point out that a treatment assignment with a sufficiently small Mahalanobis distance can be achieved by optimizing a quadratic function of linearly constrained binary variables. Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to accelerate rerandomization in various randomized experiments. Moreover, we derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR. Our simulation studies and real data analysis illustrate that VNSRR significantly improves the computational efficiency of rerandomization while maintaining desired statistical properties. It can sample thousands of treatment assignments within seconds, even in cases where existing methods require hours to complete the task.
2 Related Work
2.1 The Neyman-Rubin Potential Outcome Framework
@@
This study adopts the Neyman-Rubin potential outcome framework (Neyman, 1923; Rubin, 1974). We consider an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group. We denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators. $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise. Each unit $i$ has two potential outcomes, $(Y_{i}(1),Y_{i}(0))$, which correspond to the two treatment arms. The individual treatment effect for unit $i$ is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$. And the average treatment effect is defined as $\tau=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}(1)-Y_{i}(0))$. The observed outcome for unit $i$ is $Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0)$. We adopt the stable unit treatment value assumption(SUTVA) (Rubin, 1980), which states that the treatment levels are well-defined and there is no interference between $(Y_{i}(1),Y_{i}(0))$ and $\{W_{j}:j\neq i\}$. We estimate the treatment effects with the difference-in-means estimator, defined as $\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0)$. We assume that each unit has $p$ observed covariates, denoted by $X_{i}=(X_{i1},\ldots,X_{ip})^{T}$. The covariates of all units are gathered into a matrix $X=(X_{1},\ldots,X_{n})^{T}$. The corresponding covariance matrix is denoted by $S_{XX}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{T}$.
2.2 Rerandomization Using the Mahalanobis Distance
For a given assignment $W$, the difference in the covariate means of the two treatment groups is
$\overline{X}_{t}-\overline{X}_{c}$ $=\frac{1}{n_{t}}X^{T}W-\frac{1}{n_{c}}X^{T}(1_{n}-W)$
$=\frac{n}{n_{t}n_{c}}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
And the Mahalanobis distance between the covariates means is defined as
$M(W)$ $:=\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}\left[Cov[\left(\overline{X}_{t}-\overline{X}_{c}\right)]\right]^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
$=\frac{n_{t}n_{c}}{n}\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}S_{XX}^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
$=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n})$
$=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
@@ 2.2 Rerandomization Using the Mahalanobis Distance
Morgan and Rubin (2012) suggested to conduct rerandomization by sampling $W$ from $\{W:\sum_{i=1}^{n}W_{i}=n_{t},W_{i}\in\{0,1\}\}$ repeatedly until $M(W)\leq a$, where $a$ is a prespecified threshold. We refer to rerandomization based on this acceptance-rejection sampling strategy as ARSRR.
2.3 Pair-switching Rerandomization (PSRR)
To avoid the computationally inefficient acceptance-rejection sampling approach and accelerate rerandomization in both nonsequential and sequential experiments, Zhu and Liu (2022) suggested to search for balanced assignments with a pair-switching strategy, i.e., PSRR. The algorithm essentially starts with a random assignment and gradually improves it by switching between a treated unit and a control unit repeatedly. Motivated by the Metropolis-Hastings algorithm, PSRR allows moving to less balanced assignments with a non-zero probability to prevent being trapped in a local region. The authors obtained the unbiasedness and a lower bound on the variance reduction of the treatment effect estimator under PSRR. However, as pointed out in Section 1 and demonstrated by the simulation studies in Section 5, PSRR is still time-consuming in many settings and is not applicable in certain structured experiments.
3 Variable Neighborhood Searching Rerandomization(VNSRR)
We introduce our variable neighborhood searching rerandomization (VNSRR) in this section.
3.1 The VNSRR Algorithm
@@
minimize $\quad M(W)=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
@@
For simplification, we denote the feasible set of (3.1) as $\mathcal{F}_{0}=\{W\in\{0,1\}^{n}:\sum_{i}W_{i}=n_{t}\}$. While rerandomization aims to find a feasible $W$ in $\mathcal{F}_{0}$ with sufficiently small $M(W)$ rather than the minimizer of $M(W)$ in $\mathcal{F}_{0}$, we can still utilize combinatorial optimization techniques to conduct rerandomization and are allowed to terminate the optimization process as long as $M(W)$ falls below a prespecified threshold $a$.
There are few studies concerning the specific optimization problem in (3.1), but some related constrained binary quadratic programming problems have received extensive attention, such as the maximum diversity problem and the heaviest k-subgraph problem. These problems share the same feasible sets as (3.1) but have different objective functions. We recommend that interested readers refer to Martí et al. (2013) for a detailed review of the literature. The authors presented extensive experiments to reveal that the best methods to solve the maximum diversity problem are local search based metaheuristics, especially variable neighborhood search (Brimberg et al., 2009; Aringhieri and Cordone, 2011) and tabu search (Palubeckis, 2007; Aringhieri and Cordone, 2011). Both methods update their solutions by exchanging coordinates of $W$ that are ones with coordinates of $W$ that are zeroes. Variable neighborhood search systematically changes the neighborhood by executing two steps iteratively. It descents to find local optimum in local search steps and perturbs to avoid being trapped in local regions in shaking steps. Tabu search maintains a list of forbidden moves, namely the tabu list, to avoid looping over recently visited coordinates of $W$. Aringhieri and Cordone (2011) proposed to incorporate tabu search in the local search step of variable neighborhood search.
@@
Set $t=0$;
Set $W^{(0)}$ as $n_{t}$ elements equal to 1 and $n_{c}$ elements equal to 0 with random positions;
Set $M^{(0)}=M(W^{(0)})$;
@@
                   Set $\widetilde{S}=0$;
                   Set $t=t+1$;
                   Set $W^{(t)}=W^{*}$;
                   Set $M^{(t)}=M^{*}$;
                   if $M^{(t)}<a$ then break;
@@
Algorithm 1 Variable neighborhood searching rerandomization
Motivated by the empirical success of local-search-based metaheuristics in related optimization problems, we propose a novel VNSRR method for rerandomization through variable neighborhood searching, as summarized by Algorithm 1. VNSRR starts with a random assignment and then executes the local search step and the shaking step iteratively until an acceptable assignment is reached. VNSRR conducts the local search step differently from any existing variable neighborhood search methods. It is inspired by but different from tabu search methods. In the local search step, VNSRR sequentially examines $L$ disjoint random pairs of treated units and control units and swaps a pair if it leads to a more balanced assignment. We borrow the idea of preventing looping over recently visited units from tabu search. However, VNSRR differs from tabu search in that VNSRR avoids maintaining the tabu list and checking whether each swap is forbidden, further improving computational efficiency. We note that when selecting multiple treated units or control units in the local search step, the ordering of the units is also randomly selected. In other words, we randomly select a permutation of $L$ positions of the 1’s or 0’s in $W^{(t)}$. In the shaking step, VNSRR generates an assignment in the neighborhood of the current assignment by randomly switching $\widetilde{S}$ pairs of treated and control units. When no pairs are switched in the local search step, indicating a likely local optimum, $\widetilde{S}$ is set as a positive value $S$ to prevent getting stuck. Otherwise, $\widetilde{S}$ is set as 0, and no perturbation is performed to allow for further search in the current local region. We compare different choices of $L$ and $S$ in our experiments, and the results are relegated to the Supplement.
We point out that both PSRR and VNSRR use pairwise swap as the basic operation in their algorithms, but use it with different strategies. PSRR allows switching recently considered units, while VNSRR does not. PSRR, as motivated by simulated annealing and Metropolis-Hastings algorithm, accepts swaps that lead to worse covariate balance, i.e., larger $M(W)$, with a certain probability determined by a temperature parameter. In contrast, VNSRR keeps moving to new $W$ with smaller $M(W)$ in the local search step until none of the $L$ swaps lead to smaller $M(W)$, and then the shaking step is conducted. Our experiments illustrate that PSRR can move to $W$ with worse $M(W)$ quite frequently, while the strategy of VNSRR leads to a faster descent of $M(W)$. In fact, PSRR with zero temperature is equivalent to VNSRR with $L=1$ and without shaking steps. Our experiments show that this special version lies in the middle of PSRR and VNSRR. The zero-temperature version of PSRR was not considered by the authors of PSRR and is faster than their default version. Our default VNSRR is faster than VNSRR with $L=1$ and without shaking steps.
Like PSRR, we adopt the following trick to update the Mahalanobis distance efficiently. When we obtain a new assignment $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$, $M(W^{*})$ can be efficiently calculated as
$M(W^{*})=$ $M^{(t)}-2\sum_{i=1}^{n}W_{i}^{(t)}H_{O_{l}i}+H_{O_{l}O_{l}}+2\sum_{i=1}^{n}W_{i}^{*}H_{Z_{l}i}-H_{Z_{l}Z_{l}}+h_{O_{l}}-h_{Z_{l}},$
where $h=2\frac{n_{t}}{n}H1_{n}$.
We note that some deterministic or semi-deterministic algorithms might generate a single acceptable assignment faster than VNSRR. However, such algorithms cannot be used to generate numerous independent assignments and do not guarantee the unbiasedness of the corresponding treatment effect estimator. The next section discusses the statistical properties of different rerandomization methods.
3.2 Statistical Properties of VNSRR
In this section, we establish the statistical properties of various rerandomization algorithms based on the Mahalanobis distance. For clarification, a rerandomization algorithm here is a sampling algorithm that randomly generates a balanced treatment assignment from the feasible set $\mathcal{F}\subseteq\mathcal{F}_{0}$ by generating a chain of treatment assignments, $(W^{(0)},\ldots,W^{(T)})$, where $W^{(t)}\in\mathcal{F}$ for all $t$ and $W^{(T)}\in\mathcal{F}\cap\{W:M(W)\leq a\}$ is the final output. We provide sufficient conditions for a rerandomization algorithm to yield unbiased treatment effect estimators with bounded variance. Specifically, we consider the following conditions, which assumes Morgan and Rubin (2012) as well as Zhu and Liu (2022).
Condition 1

### v2

# Fast Rerandomization via the BRAIN

@@ Abstract
Randomized experiments are a crucial tool for causal inference in many different fields. Rerandomization addresses any covariate imbalance in such experiments by resampling treatment assignments until certain balance criteria are satisfied. However, rerandomization based on naïve acceptance-rejection sampling is computationally inefficient, especially when numerous independent assignments are required to perform randomization-based statistical inference. Existing acceleration methods are suboptimal and not applicable in structured experiments, including stratified and clustered experiments. Based on metaheuristics in integer programming, we propose BRAIN — a novel computationally-lightweight methodology that ensures covariate balance in randomized experiments while significantly accelerating the computation. Our BRAIN method provides unbiased treatment effect estimators with reduced variance compared to complete randomization, preserving the desirable statistical properties of traditional rerandomization. Simulation studies and a real data example demonstrate the benefits of our method in fast sampling while retaining the appealing statistical guarantees.
@@ 1 Introduction
Randomized experiments have been the gold standard for drawing causal inferences in a wide range of settings, including clinical trials [33], social sciences [6], and economic research [14]. The method relies on the assumption that, by randomly assigning units to treatment and control groups, covariates are balanced. Therefore, any differences in outcomes between the groups can be attributed causally to the treatment. However, covariate imbalances between treatment groups often exist in practice due to the randomness in such complete randomization (CR) [34, 42, 29, 21]. This is a major issue because it introduces confounding, making it unclear whether observed differences in outcomes are due to the treatment or to pre-existing differences between groups, consequently invalidating any inference made.
@@
An intuitive solution to this issue is rerandomization – randomizing repeatedly until some prespecified covariate balance criterion is achieved [39, 12, 3, 26, 19, 10]. The most popular choice for the balance criterion is requiring the Mahalanobis distance between the covariate matrices of the treatment group and the control group to be lower than some threshold $a$. Despite the long history and widespread use of rerandomization, its theoretical implications using the Mahalanobis distance was only recently studied by [29]. Since then, rerandomization has attracted growing research interests, and its theory underpinnings have been established in various scenarios [30, 22, 45, 23, 38, 40, 24, 44, 8].
Classical rerandomization methods repeatedly sample treatment assignments until a balanced one is generated. However, in randomization-based inference, we need to generate a large number of balanced treatment assignments to ensure reliable estimation and inference. This often requires performing thousands of rerandomizations to obtain sufficiently many independent balanced assignments. As a consequence, rerandomization is extremely time-consuming using the acceptance-rejection sampling rerandomization (ARSRR) strategy [25, 11, 5]. This pose practical challenges, especially in large-scale or real-time experiments where treatment assignments must be generated quickly. High computational cost may limit the number of balanced assignments available for randomization-based inference, and thus discourage practitioners from adopting rerandomization methods altogether.
Several methods have been proposed to alleviate the computational burden of rerandomization. However, these approaches often fall short in practice, particularly in challenging settings involving diminishing covariate imbalance or high-dimensional covariates [7, 41, 43]. Meanwhile, some alternative approaches — such as deterministic or semi-deterministic algorithms — may yield a single acceptable assignment more quickly. Nevertheless, these methods typically lack the ability to generate independent assignments and do not ensure the unbiasedness of the resulting treatment effect estimator [15, 33, 20].
To address these limitations, [46] introduced the Pair-Switching Rerandomization (PSRR) algorithm to accelerate the search for balanced assignments in both nonsequential and sequential experiments. The algorithm starts with a random assignment and gradually improves it by switching between a treated unit and a control unit repeatedly. Motivated by the Metropolis-Hastings algorithm [28, 17], PSRR allows moving to less balanced assignments with non-zero probabilities determined by a temperature hyperparameter to prevent being trapped in a local region. While the authors establish the unbiasedness and a lower bound on the variance reduction of the treatment effect estimator under PSRR, the PSRR method remains computationally intensive in many scenarios as shown in Section 5.
Structured experiments are also common in practice. The ability to adapt to these special structures are also important for rerandomization algorithms. Rerandomization methods have been proposed for sequential experiments where units arrive group by group [45, 4]; stratified experiment where experimenters divide units into strata and conduct randomization within each stratum [16, 40]; and clustered experiments where all units in the same cluster receives the same treatment [24]. The applicability of PSRR is limited to sequential experiments only. Beside CR and the naïve ARSRR, there does not exist any fast and unified framework that works in all these structured settings. Therefore, we aim to propose a method that is fast and adaptable to all structured settings.
In this paper, we novelly consider reformulating the problem from a constrained optimization (rather than sampling) perspective. For an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group; we denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators, where $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise. For some covariate balance criterion $M(W)$ — for instance — the popular choice of Mahalanobis distance [29], we consider the optimization problem
minimize $\quad M(W),$
subject to $\quad W\in\Big{\{}W\in\{0,1\}^{n}:\sum\nolimits_{i}W_{i}=n_{t}\Big{\}}.$ (1)
Under this formulation, we consider utilizing combinatorial optimization techniques to conduct rerandomization with early termination. There are few studies concerning this specific optimization problem, but some related constrained binary quadratic programming problems have received extensive attention, such as the maximum diversity problem and the heaviest k-subgraph problem. These problems share the same feasible sets as (1) but have different objective functions. We recommend interested readers refer to [27] for a detailed review of the literature. The authors presented extensive experiments to reveal that the best methods to solve the maximum diversity problem are local search based metaheuristics, especially variable neighborhood search [9, 1] and tabu search [32, 1]. Both methods update their solutions by exchanging coordinates of $W$ that are ones with coordinates of $W$ that are zeroes. Variable neighborhood searching systematically changes the neighborhood by executing two steps iteratively. It descents to find local optimum in local search steps and perturbs to avoid being trapped in local regions in shaking steps. Tabu search maintains a list of forbidden moves, namely the tabu list, to avoid looping over recently visited coordinates of $W$. Furthermore, [1] proposed to incorporate tabu search in the local search step of variable neighborhood search.
@@
Following these intuitions, we propose a novel method to accelerate rerandomization — Balanced RAdomization via INteger programming (BRAIN). We show that a treatment assignment with a sufficiently small Mahalanobis distance can be achieved by optimizing a quadratic function of linearly constrained binary variables. Moreover, we prove that BRAIN, as a member of the rerandomization (RR) family, leads to treatment effect estimators that are unbiased and has lower variances than CR. Our simulation studies and real data analysis illustrate that BRAIN significantly improves the computational efficiency of rerandomization while maintaining the desired statistical properties. Using our method, the time it takes to obtain thousands of balanced treatment assignments can be reduced to seconds when existing methods require hours.
2 Problem Setup
In this section, we formally formulate the rerandomization framework and discuss existing approaches to solve the problem.
@@ 2 Problem Setup
The Neyman-Rubin Framework. We adopt the Neyman-Rubin potential outcome framework [31, 35]. We consider an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group. We denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators, where $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise. We assume that each unit has $p$-dimensional covariates, denoted by $X_{i}=(X_{i1},\ldots,X_{ip})^{T}$. We thus denote the covariate matrix as $X=(X_{1},\ldots,X_{n})^{T}$, with corresponding covariance matrix $S_{XX}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{T}$. We assume that each unit $i$ has two potential outcomes, $(Y_{i}(1),Y_{i}(0))$, which correspond to the two treatment arms.
We denote the individual treatment effect for unit $i$ as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and the average treatment effect as $\tau=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}(1)-Y_{i}(0))$. The observed outcome for unit $i$ is $Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0)$. We adopt the stable unit treatment value assumption (SUTVA) [36], which states that the treatment levels are well-defined, and there is no interference between the potential outcomes $(Y_{i}(1),Y_{i}(0))$ and the treatment assignments of other units $\{W_{j}:j\neq i\}$. We estimate the treatment effects with the difference-in-means estimator, defined as
$\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0).$
We employ randomization-based inference to construct confidence intervals for $\tau$ and test for the sharp null hypothesis $H_{0}:\tau_{i}=0$, $i=1,\ldots,n$. See Appendix C for an introduction to randomization-based inference, including the Fisher randomization test. The procedure requires generating sufficiently many (e.g., 1000) i.i.d. assignments.
Rerandomization with the Mahalanobis Distance. [29] suggested to conduct rerandomization by sampling $W$ from $\{W:\sum_{i=1}^{n}W_{i}=n_{t},W_{i}\in\{0,1\}\}$ repeatedly until $M(W)\leq a$, where $a$ is a prespecified threshold and $M(W)$ is the Mahalanobis distance between the covariates matrices of the treatment groups. $M(W)$ is formally defined as
$M(W)$ $:=\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}\left[\textnormal{Cov}[\left(\overline{X}_{t}-\overline{X}_{c}\right)]\right]^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
$=\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}H\left(W-\frac{n_{t}}{n}1_{n}\right),$
@@
Though this approach is commonly used in practice, it is computationally expensive. We will now introduce our BRAIN method, which addresses the computational challenge by leveraging metaheuristics from integer programming.
3 The BRAIN Method
In this section, we first introduce our Balanced RAdomization via INteger programming (BRAIN) algorithm for rerandomization. We then establish statistical properties of the resulting difference-in-means estimator using BRAIN.
3.1 The BRAIN Algorithm
@@
minimize $\quad M(W)=\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}H\left(W-\frac{n_{t}}{n}1_{n}\right),$
@@
For simplification, we denote the feasible set of (3.1) as $\mathcal{F}_{0}=\{W\in\{0,1\}^{n}:\sum_{i}W_{i}=n_{t}\}$. While rerandomization aims to find a feasible $W$ in $\mathcal{F}_{0}$ with sufficiently small $M(W)$ rather than the minimizer of $M(W)$ in $\mathcal{F}_{0}$, we can still utilize combinatorial optimization techniques to conduct rerandomization with early termination (as long as $M(W)$ falls below a prespecified threshold $a$).
Motivated by the empirical success of local-search-based metaheuristics in related optimization problems, we propose the BRAIN method as follows. BRAIN starts with a random treatment assignment. Then, it conducts a local search step where we sequentially examine $L$ disjoint random pairs of treated units and control units and swaps a pair if it leads to a more balanced assignment. If any swap has been made in the local search step, BRAIN will continue to search in the current local region. When no pairs are switched in the local search step, indicating a likely local optimum, BRAIN performs a perturbation step by randomly switching $S$ pairs of treated and control units. The algorithm will repeat the local search step and the perturbation step until reaching a satisfiable sample with $M(W)\leq\alpha$. The full algorithm is shown in Algorithm 1.
BRAIN conducts the local search step differently from any existing variable neighborhood searching methods. We borrow the idea of preventing looping over recently visited units from tabu search and examine $L$ disjoint pairs. However, BRAIN differs from tabu search in the sense that BRAIN avoids maintaining the tabu list and checking whether each swap is forbidden, further improving computational efficiency.
$L$ and $S$ are two hyperparameters in BRAIN. If $L$ is too small, the algorithm would still loop over recently visited pairs. If $S$ is too large, the perturbation is too strong such that it is similar to random restart. We compare different choices of $L$ and $S$ in Appendix D and suggest setting $(L,S)=(n/2,1)$.
Our method differs from PSRR in that it accepts swaps only if $M(W)$ decreases in local search step. This way, it avoids inefficiently moving to a treatment assignment $W$ with higher $M(W)$, which frequently happens for PSRR. By avoiding this issue, our algorithm leads to significant faster descent of $M(W)$ and therefore much shorter computation time. PSRR with zero temperature is equivalent to BRAIN with $L=1$ and without the perturbation steps. Our experiments show that this special version of PSRR is faster than its default version but still much slower than BRAIN.
@@
Set $t=0$, $W^{(0)}$ randomly with $n_{t}$ many 1’s and $n_{c}$ many 0’s, $M^{(0)}=M(W^{(0)})$;
@@
                   Set $\widetilde{S}=0$, $t=t+1$, $W^{(t)}=W^{*}$, $M^{(t)}=M^{*}$;
                   if $M^{(t)}\leq a$ then break;
             end if
       end for
@@
Algorithm 1 Balanced RAdomization via INteger programming (BRAIN)
3.2 Statistical Properties of BRAIN
In this section, we establish the statistical properties of BRAIN. We consider the following conditions, which are standard in the literature [29, 46].
Condition 1.